OpenStudy (anonymous):

Which ordered pair is a solution of y=2x+1

sam (.sam.):

Ordered pair?

OpenStudy (anonymous):

the ordered pairs that are available to chose are (3,7) (0,1) (-1,1) (-3,-5)

sam (.sam.):

Well from here y=2x+1 If you substitute both values x and y from (x,y), 7=2(3)+1 7=6+1 7=7 Oh bingo! (3,7) is the solution